In an air compressor air flows steadily at the rate of 0.5 kg/s through an

air compressor. It enters the compressor at 6 m/s with a pressure of 1 bar and a specific volume

of 0.85 m3/kg and leaves at 5 m/s with a pressure of 7 bar and a specific volume of 0.16 m3/kg. The

internal energy of the air leaving is 90 kJ/kg greater than that of the air entering. Cooling water

in a jacket surrounding the cylinder absorbs heat from the air at the rate of 60 kJ/s. Calculate :

(i) The power required to drive the compressor ;

(ii) The inlet and output pipe cross-sectional areas



1
Expert's answer
2021-08-26T14:48:08-0400


Mass flow rate of air m = 0.5 kg/s

Velocity at inlet "V_1 = 6 \; m/s"

Pressure "P_1 = 1\; bar = 10^5 \;Pa"

Specific volume "v_1 = 0.85 \;m^3/kg"

Outlet velocity "V_2 = 5 \;m/s"

Pressure "P_2 = 7 \;bar = 7 \times 10^5 \;Pa"

Specific volume "v_2 = 0.16 \;m^3/kg"

Internal energy U = 90 kJ/kg

Heat absorbed Q = -60 kJ/s

Power required W =?

"∆Pv = P_2v_2 – P_1v_1 \\

= 7 \times 10^5 \times 0.16 - 10^5 \times 0.85 \\

= 27000 \; J/kg \\

= 27 \; kJ/kg \\

∆h = ∆(U + Pv) \\

= 90 + 27 \\

= 117 \; kJ/kg"

Change in kinetic energy

"∆KE = 0.5(V_2^2 - V_1^2) \\

= 0.5(5^2 - 6^2) \\

= - 5.5 \; J/kg"

Energy balance equation

"Q - W = m(∆h + ∆KE) \\

-60 - W = 0.5(117 - \frac{5.5}{1000}) \\

W = - 118.5 \; kW"

Negative sign shows that work is done on compressor.

Inlet and outlet pipe cross-sectional areas, "A_{1}" and "A_{2}" :

Using the relation,

"m = \frac{VA}{v} \\

A_1 = \frac{mv_1}{V_1} = \frac{0.5 \times 0.85 }{6}=0.0708 \;m^2 \\

A_2 = \frac{mv_2}{V_2}= \frac{0.5 \times 0.16}{5}= 0.016 \;m^2"